# HouseLLM: LLM-Assisted Two-Phase Text-to-Floorplan Generation

### v3

@@ 3.2.2 Network
Discrete Decoder. Our approach aims to capture geometric relationships (such as collinearity, orthogonality, or corner adjacency), which are difficult to maintain through continuous coordinate regression alone. As in [23], we infer coordinates in discrete form. After obtaining $C^{t-1}_{i,j}$ and $C^{0}_{i,j}$ from the denoising process, we map $C^{0}_{i,j}$ back to the range [0,255] through an affine transformation, round the values, and convert them to binary representation using an “int2bit” function, resulting in an 8-dimensional binary vector. For each corner, we apply a linear transformation to generate a 512-dimensional embedding vector:
$\mbox{Linear}([\mbox{AU}(C^{g}_{i,j})_{t},\mbox{AU}(C^{l}_{i,j})_{t},R_{i},\mathbf{1}(i),\mathbf{1}(j),t])$ (9)
@@ 3.2.3 Training loss
We train our model with the simple L2-norm regression loss by [7]. The loss function consists of noise prediction loss and corner coordinate regression loss. Specifically, $\epsilon_{\theta}(C_{i,j},t)$ is compared with the ground-truth noise to quantify the discrepancy between them. The discrepancy is typically measured using the masked mean squared error as follows:
$Loss_{n}=\dfrac{\sum_{i=1}^{N}\|\epsilon_{\theta}(C_{i,j},t)-\epsilon\|_{2}}{N}$ (10)
where $N$ represents total number of samples. $C_{i,j}$ is compared with the ground-truth corner coordinate in binary representation.
$Loss_{r}=\dfrac{\sum_{i=1}^{N}\|C^{s}_{i,j}-C^{t}_{i,j}\|_{2}}{N}$ (11)
@@ 4 Experiments
Table 1: Generation performance of different methods.
Method Realism ($\uparrow$) Diversity ($\downarrow$) Compatibility ($\downarrow$)
6 5 6 7 8 5 6 7 8
House-GAN[18] -0.95 37.5 41.0 32.9 66.4 2.5 2.4 3.2 5.3
House-GAN++[19] -0.52 30.4 37.6 27.3 32.9 1.9 2.2 2.4 3.9
HouseDiffusion[23] -0.19 11.2 10.3 10.4 9.5 1.5 1.2 1.7 2.5
PuzzleFusion[9] - 10.55 0.97
Ours -0.03 8.6 7.5 8.1 9.0 0.24 0.25 0.28 0.32
Figure 5: Generation samples from HouseLLM and HouseDiffusion. The results of HouseLLM align well with user requirements (as translated from the of original data samples) in the number and type of rooms, also in the overall room distribution. HouseDiffusion generally performs well, but in some cases, it can generate gaps or holes within the house (indicated by the dashed-line rectangles), along with misaligned object placement, which make the layout unrealistic.
@@ 4.1 Datasets and implementation details
Our method is implemented in PyTorch and and accelerated using an NVIDIA Quadro RTX 8000. The batch size is set to 128 and learning rate is 1e-3. Our models are optimized by Adam [11] with weight decay [12] for 250K steps and the initial learning rate is 1e-3. The network architecture can be found in Appendix.
Datasets. We use the commonly used public dataset RPlan [32] for experiments. RPlan represents the largest dataset for house plan with over 80K images. We divide RPlan into five groups (5, 6, 7, 8 rooms) according to the number of rooms for cross validation. Ablation experiments were conducted on the 6-rooms. To create paired groups for training, we extracted 10K initial solutions from the LLM using the prompting method from Section 3.1. GPT-4o was used as the LLM in our approach.
Metrics. Following the evaluation method used in the prior work [18], we employ Realism, FID and Compatibility as evaluation metrics. Realism is an estimate based on a user survey, considering the user’s subjective experience. Diversity is measured by the FID score, defined as the Fréchet Inception Distance[6] between the two Gaussian distributions. It is used to measure the similarity between the generated image and the real image. The Compatibility score[1] is based on a graph editing distance between the layout sample and the Layout-Final. It is used to measure the distance between generated layout and the ground truth.
@@ 4.2 Experimental results
Table 1 shows the main results. For the existing methods, we reproduced the results reported in their respective papers. It can be seen that our method achieves superior performance in all metrics. For instance, compared with HouseDiffusion, HouseLLM brings an average improvement of 28% in diversity and 79% in compatibility. The significant improvements indicate that it better meets user needs in terms of flexibility in layout generation and spatial utilization efficiency.
For the realism score, we followed the same process as House-GAN++ to obtain the score. We created 1000 floorplan samples for our method and showed participants two samples from different methods, asking them to select ”A is better,” ”B is better,” or ”Both are equally good.” Method A is awarded +1, -1, or 0 points according to the participants’ choices. Each pair of methods is evaluated 150 times by 10 participants (i.e., 15 evaluations per participant). The survey results showed that 65% of participants believed that our results were as good as or better than ground truth. Table 1 shows that HouseLLM achieves significant improvements compared to other methods, with an excellent score of -0.03, suggesting that the house layouts generated by our method are visually close to realistic layouts and meet user expectations.
Figure 5 presents a visual comparison of the outputs from HouseLLM and HouseDiffusion. HouseLLM generates results that well align with the original data samples in the number and type of rooms. While HouseDiffusion generally performs well, it can generate layouts with misalignment and incorrect object placement. The primary reason is that with HouseDiffusion, the generation conditions are applied only during the denoising stage, restricting the model’s ability to accurately realize the spatial relationships between rooms throughout the generation (see more details in our ablation study). Furthermore, while the bubble diagram based conditions convey basic connectivity between rooms, they lack the granularity necessary to guide detailed spatial arrangements, such as room size and precise positioning. In contrast, HouseLLM combines initial layouts generated by LLM with a diffusion model that applies conditions during both noise addition and denoising stages, producing more practical house layouts. More implementation details can be found in Appendix.
@@ 4.3 Ablation study
Alternative generation methods. We compare our method with two alternative methods: the one-phase Test-to-Layout method that maps text directly to layout, and using LLM only to generate the results, that is, Layout-LLM. Figure 6 presents the comparison results of these methods under the same generation conditions. From the figure, we can see that the results differ greatly, and with the one-phase method, text description is not enough to support the model to maintain sufficient geometric constraints. In addition, we evaluate the primary quantitative results for the one-phase and two-Phase methods. Table 2 presents the main quantitative evaluation. Diversity indicates that the distribution of Text-to-Layout method result differs significantly from RPlan. This is due to a mismatch between the text dataset complexity and task, requiring extensive data support. This limitation restricts the model’s generalization ability, preventing it from effectively learning and capturing the true distribution of RPlan. The details of Text-to-Layout are in Appendix.
When comparing the Layout-LLM only method and the full version of our method, we can observe that former produces many overlapping rooms and fails to maintain a reasonable distribution; See Figure 6 (c) and (d). These limitations suggest that while the LLM-generated initial layout can capture the general structure of a house layout, it cannot enforce fine-grained spatial constraints. This is where the need for a refinement step becomes essential.
Figure 6: Examples of house layouts generated by three methods: the one-phase method that maps text directly to layout, LLM only, and our two-phase method.
Table 2: Comparison between the one-phase and our two-phase methods.
Method Diversity($\downarrow$) Compatibility($\downarrow$)
One phase: Text-to-Layout 109.5 10.6
Two phase: HouseLLM 7.5 0.25
Table 3: Effect of rate of condition.
Rate Diversity ($\downarrow$) Compatibility ($\downarrow$)
2e-1 8.16 25.67e-2
1e-1 8.19 24.95e-2
1e-2 8.24 23.46e-2
1e-3 7.47 25.44e-2
Table 4: Effect of forward v.s. reverse conditioning.
Forward Reverse Divers.($\downarrow$) Compat.($\downarrow$)
@@
Rate of Condition. We investigated the impact of different conditional participation ratios on the model’s generation results, as shown in Table 3. The results indicate that as the conditional ratio increases, the model performance gradually decreases, suggesting that an excessive amount of conditional information causes the model to deviate from the true distribution during generation. This may be due to the high ratio of conditional guidance, which makes the model overly dependent on the conditional information, thus weakening its ability to learn the original data distribution. Therefore, properly controlling the conditional ratio can maintain the accuracy of generated samples.
Forward v.s. Reverse Conditioning Evaluation. Different from the diffusion model used in HouseDiffusion, we incorporated conditional constraints in both the forward and reverse processes in our design. To assess this approach, we analyzed the impact of each constraint by varying their proportions independently. As shown in Table 4, adding conditional information during the noise addition phase yields better results than adding it only during the denoising phase, as it enables the model to build condition-related features earlier. Single-stage conditioning approach may lead to weaker constraints, resulting in inconsistent room sizes and irregular spatial relationships. The best performance is achieved when conditions are added in both phases, validating the effectiveness of the dual-phase approach. This strategy significantly enhances both accuracy and consistency, resulting in more precise and aligned outputs.
@@ 5 Conclusion
This paper proposes a two-Phase Text-to-Floorplan generation model, enabling user-friendly house layout generation. This work guides the reasoning of the LLM through chain-of-thought prompting to generate an initial layout based on user requirements. The initial layout is then refined using a diffusion model to produce the final house layout. Experimental results show that our method achieves state-of-the-art performance. Given the potential of LLMs, we expect that our solution be extended to the generation of other complex architectural designs such as shopping malls and office buildings. This will be a subject of our future study.

### v4

@@ 3.2.2 Network Architecture
Discrete Decoder. Our approach aims to capture geometric relationships (such as collinearity, orthogonality, or corner adjacency), which are difficult to maintain through continuous coordinate regression alone. As in Shabani et al. (2023), we infer coordinates in discrete form. After obtaining $C^{t-1}_{i,j}$ and $C^{0}_{i,j}$ from the denoising process, we map $C^{0}_{i,j}$ back to the range [0,255] through an affine transformation, round the values, and convert them to binary representation using an “int2bit” function, resulting in an 8-dimensional binary vector. For each corner, we apply a linear transformation to generate a 512-dimensional embedding vector:
$\mbox{Linear}([\mbox{AU}(C^{0}_{i,j}),\mbox{AU}(C^{0}_{i,j})_{I},\mbox{int2bit}(C^{0}_{i,j}),R_{i},\mathbf{1}(i),\mathbf{1}(j),t]).$ (9)
@@
Table 1: Generation performance of different methods.
Method Realism $\uparrow$ Diversity $\downarrow$ Compatibility $\downarrow$
Task 6 5 6 7 8 5 6 7 8
House-GANNauata et al. (2020) -0.95 37.5 41.0 32.9 66.4 2.5 2.4 3.2 5.3
House-GAN++Nauata et al. (2021) -0.52 30.4 37.6 27.3 32.9 1.9 2.2 2.4 3.9
HouseDiffusionShabani et al. (2023) -0.19 11.2 10.3 10.4 9.5 1.5 1.2 1.7 2.5
PuzzleFusionHossieni et al. (2024) - 10.55 0.97
Tell2Design Leng et al. (2023) -1.03 42.74 -
Ours -0.03 8.6 7.5 8.1 9.0 0.24 0.25 0.28 0.32
@@ 3.2.3 Training Loss
We train our model with the simple L2-norm regression loss by Ho et al. (2020). The loss function consists of noise prediction loss and corner coordinate regression loss. Specifically, $\epsilon_{\theta}(C_{i,j},t)$ is compared with the ground-truth noise to quantify the discrepancy between them. The discrepancy is typically measured using the masked mean squared error as follows:
$Loss_{n}=\dfrac{\sum_{i=1}^{N}\|\epsilon_{\theta}(C_{i,j},t)-\epsilon\|_{2}}{N},$ (10)
where $N$ represents total number of samples. $C_{i,j}$ is compared with the ground-truth corner coordinate in binary representation:
$Loss_{r}=\dfrac{\sum_{i=1}^{N}\|C^{s}_{i,j}-C^{t}_{i,j}\|_{2}}{N}.$ (11)
@@ 4.1 Datasets and implementation details
Our method is implemented in PyTorch and runs on an NVIDIA Quadro RTX 8000. The batch size is set to 128 and the learning rate is 1e-3. Our models are optimized by Adam Kingma (2014) with weight decay Loshchilov (2017) for 250K steps, with the initial learning rate is 1e-3. The network architecture can be found in the Appendix.
Datasets. We use the public dataset RPlan Wu et al. (2019) for experiments. RPlan represents the largest dataset for floorplan with over 80K images. We divide RPlan into five groups (5, 6, 7, 8 rooms) according to the number of rooms for cross-validation. Ablation experiments were conducted on the 6-room task. To create paired groups for training, we extracted 10K initial solutions from the LLM using the prompting method from Section 3.1. GPT-4o was used as the LLM in our approach.
Metrics. Following prior work Nauata et al. (2020), we employ Realism, FID, and Compatibility as evaluation metrics. Realism is an estimate based on a user survey, considering the user’s subjective experience. Diversity is measured by the FID score, defined as the Fréchet Inception DistanceHeusel et al. (2017) between the two Gaussian distributions. It is used to measure the similarity between the generated images and the real image. The Compatibility scoreAbu-Aisheh et al. (2015) is based on a graph editing distance between the layout sample and the Layout-Final. It is used to measure the distance between the generated layout and the ground truth. In addition, Macro IoU and Micro IoU are introduced for comparison with the Text-to-Layout method Tell2Design Leng et al. (2023). Macro IoU computes the average IoU across different room types, whereas Micro IoU calculates the global IoU by aggregating the IoUs of all rooms.
@@ 4.2 Experimental results
Figure 5: Generation samples from Tell2Design, HouseDiffusion and HouseTune. The results of HouseTune align well with user requirements in terms of room count and type, with plausible room arrangement. The results produced by Tell2Design exhibit high similarity to the Reference layouts, demonstrating limited diversity. However, the number of generated houses deviates from the Reference, indicating inconsistencies in quantity. HouseDiffusion performs reasonably well; yet, in some cases, it generates gaps or holes within the house (indicated by the dashed-rectangles), along with misaligned object placement, which makes the layout unrealistic.
Table 1 presents the main results, where we reproduce the reported performance of existing methods for fair comparison. As shown in the table, our approach consistently outperforms prior methods across all evaluation metrics. Specifically, compared to HouseDiffusion, our method improves diversity by 28% and compatibility by 79%, demonstrating its effectiveness in generating flexible and spatially efficient layouts. Tell2Design is evaluated using the IoU metric, with 54.34% Micro IoU and 53.30% Macro IoU. In comparison, HouseTune, when using unlabeled data, achieves a Micro IoU of 21.84% and a Macro IoU of 17.75%, values that are significantly lower than the results of Tell2Design with labeled data. However, due to fundamental differences in task objectives, a direct comparison based solely on IoU would be unfair. Tell2Design employs an autoregressive Seq2Seq framework, in which generation is strictly dependent on sequence order. This results in high IoU scores but significantly limits diversity (see Fig.5(B)), as it demands a large amount of labeled data to capture layout variations. In contrast, our non-autoregressive approach inherently mitigates these limitations, leading to a 79.88% improvement in diversity compared to Tell2Design. Furthermore, Tell2Design relies on a costly training process involving artificial pre-training followed by fine-tuning on human annotations. In contrast, our LLM-driven method eliminates the need for manual annotation while still achieving a higher user absolute preference score (-0.03 v.s. -1.03). Notably, under the “Training on Artificial Instructions Only" setting—comparable to our setup—HouseTune surpasses Tell2Design, with a Micro IoU of 15.69, compared to 9.13 for Tell2Design (+71.85%), and a Macro IoU of 11.43, compared to 6.06 (+88.61%). For realism evaluation, we follow the same procedure as House-GAN++ to ensure comparability. User surveys indicate that 65% of participants perceive our generated layouts as comparable to the ground truth. Implementation details are in the Appendix.
Figure 5 presents a visual comparison of layouts generated by HouseTune, Tell2Design, and HouseDiffusion. The layout descriptions for HouseTune, Tell2Design’s annotations, and HouseDiffusion’s graphs are all derived from the reference samples shown in Fig.5(a). Due to the limitation of the plain Seq2Seq model, Tell2Design often produces inaccurate room counts, resulting in unrealistic layouts (see Fig.5(B)). In contrast, HouseTune and HouseDiffusion explicitly enforce these constraints, producing more structured and diverse layouts. HouseDiffusion, despite generally producing reasonable layouts, often exhibits misaligned and misplaced objects. This issue stems from its conditioning strategy, which applies constraints only during the denoising phase, making it less effective in maintaining spatial consistency throughout the generation process (see our ablation study for further analysis). By contrast, HouseTune refines an initial LLM-generated layout using a diffusion model with conditioning applied in both the noise addition and denoising stages. This dual-phase conditioning enhances structural coherence and spatial accuracy, leading to more practical and well-organized house layouts.
@@ 4.3 Ablation study
One-stage v.s. two-stage Methods. We compare the one-stage Text-to-Layout approach using the text from LLM with our two-stage method. Figure 6 presents the comparison results under identical generation conditions. As shown in the figure, the one-stage method frequently produces overlapping rooms and fails to maintain a reasonable spatial distribution. This limitation arises because textual descriptions alone do not provide sufficient geometric constraints that enable the LLM to generate well-structured layouts. Additionally, we conduct a quantitative evaluation of both the one-stage and two-stage methods. Table 2 presents the main quantitative evaluation results. Diversity indicates that the distribution of Text-to-Layout method result differs significantly from RPlan. This is due to a mismatch between the text dataset complexity and task, requiring extensive data support. This limitation restricts the model’s generalization ability, preventing it from effectively learning and capturing the true distribution of RPlan. The details of Text-to-Layout are in the Appendix.
Table 2: Comparison between one-stage and two-stage methods.
Method Diversity $\downarrow$ Compatibility $\downarrow$
1-stage: Text-to-Layout 109.5 10.6
2-stage: HouseTune 7.5 0.25
Figure 6: Effect of two stages in house layouts generation. (a) reference; (b) the one-stage method that maps text directly to layout; (c) LLM only, and (d) our method.
When comparing the Layout-Init and Layout-Final, we can observe that former produces many overlapping rooms and fails to maintain a reasonable distribution; See Figure 6. These limitations suggest that while the LLM-generated initial layout can capture the general structure of a house layout, it cannot enforce fine-grained spatial constraints. This is where the need for a refinement step becomes essential.
Table 3: Conditional diffusion ablation study.
(a) Effect of forward v.s. reverse conditioning.
[…]
@@
(b) Effect of different conditions on the generated results.
Rate Macro IoU $\uparrow$ Micro IoU $\uparrow$
1e-1 21.37% 18.46%
1e-2 20.12% 16.59%
1e-3 18.67% 15.43%
Figure 7: Generation results with different LLMs.
Robustness and effectiveness of the prompts. Our method maintains consistent quality across both commercial and open-source LLMs, namely DeepSeek-R1 and Doubao-1.5Pro. As shown in Fig.7 and Table.4(a), the generated JSON representations preserve structural integrity across different models, enabling seamless downstream diffusion refinement without additional fine-tuning. This demonstrates that our prompting strategy exhibits low dependency on proprietary models. Furthermore, Table 4(b) presents the ablation results for different prompt designs. P1 employs only the task definition prompt to generate semantic representations. P2 incorporates a common-sense prompt, improving reasoning accuracy. P3 adopts a basic Chain-of-Thought (CoT) prompt (“Let’s think step by step” Kojima et al. (2022)), further enhancing performance. P4 introduces explicit reasoning chains, leading to additional improvements over P3. These results indicate that our tailored prompting strategy effectively enhances the quality of the initial layout generation.
Conditional diffusion ablation study. Unlike the diffusion model used in HouseDiffusion, our approach integrates conditional constraints in both the forward and reverse processes. To evaluate the effectiveness of this design, we analyzed the impact of each constraint by independently varying their proportions. As shown in Table 3(a), incorporating conditional information during the noise addition phase leads to better performance compared to applying it only during the denoising phase. This is because early integration of conditions allows the model to establish condition-related features at an earlier stage. A single-stage conditioning approach may result in weaker constraints, leading to errors in room sizes and spatial relationships. The best performance is achieved when conditions are applied in both phases, confirming the effectiveness of the dual-phase conditioning strategy.
Furthermore, we examined the effect of different conditional participation ratios on generation quality, as presented in Table 3(b). The results indicate that as the conditional ratio increases, model performance gradually declines, suggesting that an excessive amount of conditional information causes the model to diverge from the true data distribution. Therefore, properly regulating the conditional ratio is crucial for maintaining the accuracy and realism of the generated samples.
Table 4: Prompt ablation study.
(a) Comparative analysis of different LLMs.
LLMs Diversity $\downarrow$ IoU $\uparrow$
DeepSeek-R1 94.3 22.3%
DouBao-1.5Pro 95.6 20.2%
GPT-4o 95.1 21.6%
(b) Comparative analysis of different prompting strategies.
Description Diversity $\downarrow$ IoU $\uparrow$
P1 Task definition prompt 98.6 12.36%
P2 Common sense prompt 92.3 14.57%
P3 Think step by step 86.5 20.69%
P4 Predefined CoT prompt 75.2 21.41%
@@ 5 Conclusion
This paper proposes a two-stage text-to-floorplan generation model, enabling user-friendly house layout generation. This work guides the reasoning of the LLM through Chain-of-Thought prompting to generate an initial layout based on user requirements. The initial layout is then refined using a diffusion model to produce the final house layout. Experimental results show that our method achieves state-of-the-art performance. Given the potential of LLMs, we expect that our solution be extended to the generation of other complex architectural designs such as shopping malls and office buildings. This will be a subject of our future study.